OpenStudy (anonymous):

Which of these ordered pairs is a solution to the inequality 3x + y > – 8 ?

OpenStudy (anonymous):

answers could be (–3, 0) (–2, –1) (0, –10) (2, –16)

OpenStudy (anonymous):

try subbing in your answer choices... let me do the first one 3x + y > -8 (-3,0) x = -3 and y = 0 now sub 3(-3) + 0 > - 8 -9 + 0 > -8 (not correct) first one is not correct, you want to try the 2nd answer choice ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

3(-2) + -1 = -7

OpenStudy (anonymous):

its not an equal sign it is >

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

-7 > -8

OpenStudy (anonymous):

is -7 greater then -8 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you just found your answer

OpenStudy (anonymous):

if your inequality comes out true, then it is a solution

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@kelliegirl33 i PMed you

OpenStudy (anonymous):

iv got another one ok?

OpenStudy (anonymous):

Which of these ordered pairs is a solution to the inequality y – 2x ≤ –3?

OpenStudy (anonymous):

need to see the ordered pairs

OpenStudy (anonymous):

@holly43 .....I answered your message

OpenStudy (anonymous):

(2, 4) (–2, 3) (3, 4) (1, –1)

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

we do this one just like the last one...by subbing y - 2x <= -3 (2,4) 4 - 2(2) <= -3 4 - 4 <= -3 0 <= -3 (not this one) (-2,3) 3 - 2(-2) <= -3 3 + 4 <= -3 7 <= -3 (nope) (3,4) 4 - 2(3) <= -3 4 - 6 <= -3 -2 <= -3 (nope) (1,-1) -1 -2(1) <= -3 -1 - 2 <= -3 -3 <= -3 (BINGO...it would have to be the last one..lol )

OpenStudy (anonymous):

thank you sooooooo much!! @kelliegirl33